OpenStudy (anonymous):

for the sequencing of 3/16, 4/25, 5/36, 6/49 would it be (n+1)/(n^2)?

OpenStudy (anonymous):

\(\large {n \over (n+1)^2}\).

OpenStudy (anonymous):

why do we square the bottom

OpenStudy (anonymous):

You see the first numbers are \(\large \frac{3}{4^2}, \frac{4}{5^2}, \frac{5}{6^2}\).

OpenStudy (anonymous):

yes i see it

OpenStudy (anonymous):

So if the numerator is \(n\), then the denominator is \((n+1)^2\).

OpenStudy (anonymous):

ohh yes now i see it clearer, you have to use the same n for the top and bottom

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

:/ i put that in my online hw and it says its incorrect....it makes sense the way you did it too

OpenStudy (anonymous):

You need to write the whole question.

OpenStudy (anonymous):

the question only asks for the formula which is n/(n+1)^2

OpenStudy (anonymous):

Why does it say wrong then? Could you write the exact problem?

OpenStudy (cwrw238):

if n means the cardinal number then the nth term you got isnt correct

OpenStudy (anonymous):

For each sequence, find a formula for the general term, \[a _{n}\]. \[1/2, 1/4, 1/6, 1/8\]. i got the answer for this one which was (1)/(n+n) and the other for the\[3/16, 4/25, 5/36, 6/49\]

OpenStudy (anonymous):

It depends on which n you start at. For the first one, you started at \(n=1\). If we do the same with the second one, we should write is as \({n+2 \over (n+3)^2}\).

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

yes i got the answer right thanks alot!!